Sol Summary – November 19th

Sol Summary Report

November 19, 2017

Seek out new worlds…

Shannon Rupert, MDRS Director

Mission Status: We are not in sim. Brandon and I are the only two
here.  Les and his team departed early.  We are awaiting Crew 183 on
Tuesday, when they will begin their one-week mission.

Sol Activity Summary:  We did some map/road checking today, exploring
road 1574 and Cactus Road.  Updates will be in my soon to be released
guide to MDRS roads.

Look Ahead Plan: Tomorrow HudgesNet comes to fix my service and we
begin work on the pad for the new RAMM, which will be installed on
campus on Tuesday.  I will also order winter blend diesel, because I
believe it is now cold enough that the summer mix is what shut down
the generator last night.  We also need to get water, ethanol free gas
and work on the new rovers.

Anomalies in work:  Not a sim day.  Generator did not run all night,
but has run all day.  I suspect the summer blend diesel is the issue.
Tried to get lower deck heater to work this morning but it did not.
Will continue troubleshooting.  Water was frozen in the am but thawed
by noon.  We need to get the tank and pipes insulated soon.  I will
contact the contractor.

Weather:  Calm, sunny and cold.

Crew Physical Status:  We are both fine.

EVA: Out of sim EVA to check out northern routes.

Reports to be file: Sol Summary and Operations Report, which was sent
earlier today.  No photos will be sent.

Support Requested: NOTICE:  Effectively immediately and until I say
differently in the spring, the heater at the Hab will stay on as

Temperature 72 degrees

Fan on auto

Heater on ON

The controller should not be adjusted at all by any crew.  When you
adjust it, we lose the heater and it just blows cold air, but if you
just leave it alone, it works perfectly.  If the crew is cold, they
should turn on the secondary heater (once I get it running!)

Please find out what type of filter the heater uses and let me know so
I can order some.  Thank you.

Operations Report – November 19th

Crew 182+ Operations Report 19NOV2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: water heater, lower deck electrical outlets, one
suit charger, generator, Hab water system, front airlock window, top

Notes: Water heater needs to be replaced. It was supposed to be
replaced two Mondays ago, but when the contractor came to do it, the
crew did not open the door

Lower deck electrical outlets: Electrician is being hired to look at this.

Suit charger: Broken by crewmember. An additional three suit chargers
were not working when the suits were delivered. Contacting NorCal
regarding replacement.

Generator failed in the night. Believed to be low oil. Will monitor.

Last night when I went into the Hab, the water pump was running
non-stop and the system was empty. After I refilled the tank to ½ and
troubleshooted, I determined that someone just ran the system dry. I
reprimed the lines and the pump began functioning normally.

The front airlock windown came out this summer, and during the refit
and again the night before last. Need to find a way to secure it.

The new top hatch is not secure in high winds. Need to find a way to secure it.

Generator— Turned off yesterday am at 8 am, turned on last night at 6
pm, failed during the night, was restarted after adding oil at 7:45
am. SOC was at 61% so I’m going to let it run all day. We need to
replace the summer blend with a winter blend and we did not use the
amount of fuel we thought we would because the generator was out of
service all summer. Note: only use the generator specific oil.

Solar— SOC was at 61% this morning. Will monitor throughout the day.

Diesel – 65%

Propane – n/a psi. Amerigas was contacted and propane was ordered.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons. Note
this should be rewritten on the template as Ethanol Free gas. This is
the only gas that should go in the ATVs. It is only available at
Stan’s Chevron. In addition, the only oil that should go in the ATVs
is an ATV specialty oil.

Water (trailer) – 0 gallons.

Water (static) – ~ 200 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: no (at least not by me)

ATV’s Used: All ATVs except the 300 were used today. All 5 have been
serviced and repaired for the season. The 300 still needs its parking
brake repaired, parts are on order. The 350s all have new seats so
none of the machines has a ripped seat. We will be using one today.

Oil Checked: no

Atv # Fuel Used Gals: Not none. We will go get gas toady.

Tires Status: Fine

# Hours the ATVs were Used today:

Deimos used: Yes

Hours: nor reported

Beginning charge:

Ending charge:

Currently charging: yes

HabCar used? Yes. We got one load of water in it today.

Notes and Comments: The four new rovers are charged but still need to
be put together and their seats modified. Progress will be reports as
I get to them.

We are using so much less water with the new toilet.

Summary of internet: Also needs to be added to template. Not sure
about Hab’s internet.

Summary of suits: This needs to be added to the template. All suits
are operational, 6 suits in EVA room. Two exploration suits in EVA
room, only one charger. Prototype exploration suit in EVA room. Two
exploration suits in Science Dome with no working chargers. I found a
bottle of soda lime on the stirrer in the lab and it was on. I turned
it off (hope it wasn’t Brandon’s, just thought about this).

Summary of Hab operations: Noted several failed systems and they are
noted above. The cement pad out the rear airlock needs some serious

Summary of GreenHab operations: Heating at 50 degrees F. It’s lovely
in there. Plants are looking good. Hose has a leak and freezes at
night. Needs to be modified to a very short hose that is taken in
every night or water taken directly from freeze free faucet. Crews
need to be sure they push the level all the way down and use it daily
to prevent freezing. Caution should be taken to not spill any water
on the ground as it will turn to dangerous ice,

Summary of ScienceDome operations: The lab is pretty much a mess.
Brandon is still doing his work, but lab will require cleaning before
next crew. Shop vac was not used nor put together.

Summary of health and safety issues: Still need alarms, will check to
be sure they were ordered.

Questions, concerns and requests to Mission Support: Please update
this template with the additions indicated. Thank you.

Copyright © The Mars Society. All rights reserved. | Main Site